how do you solve log2x=log4?
Answer by haileytucki(390)

log2x=log4
Since the logarithms on both sides of the equation have the same base, the only way the logarithms can be equal is if the arguments are equal.
2x=4
Divide each term in the equation by 2.
(2x)/(2)=(4)/(2)
Simplify the left-hand side of the equation by canceling the common factors.
x=(4)/(2)
Simplify the right-hand side of the equation by simplifying each term.
x=2